--------------- QUESTION FOLLOWS --------------------------


a company tests its workers with a lie-detector that is correct 90% of the time and fires anyone fail the test suppose 5% of the workers are guilty of theft



1- for the workers who are fired ,calculate the percentage of them are actually guilty ?



1
Expert's answer
2021-11-15T20:23:35-0500

P(Thief) = 0.05

P(Fail | Thief) = 0.9


"P(Fail) = (0.05)\times(0.9) + (0.95)\times(0.1) = 0.14"


"P(Thief | Fail) = P(Fail | Thief)\frac{P(Thief)}{P(Fail)
}"


"P(Thief | Fail) = 0.9\times\frac{0.05}{0.14}=0.3214"


This means 32.14% of them are actually guilty.